Find all zeros of P(x)= x^2+3x^2-4x-12
Answer by dabanfield(803)

Find all zeros of P(x)= x^2+3x^2-4x-12
We need to solve:
x^2 + 3x^2 - 4x - 12 = 0
Collecting terms we have;
4x^2 - 4x - 12 = 0
Divide both sides by 4:
x^2 - x - 3 = 0
Use the quadratic formula where a=1, b= -1, c= -3 to solve for x.